# Humans Hallucinate Too: Language Models Identify and Correct Subjective Annotation Errors With Label-in-a-Haystack Prompts

### v1

@@ Acknowledgments
This project was supported in part by funds from DARPA under contract HR001121C0168, NSF CIVIC, and USC-Capital One Center for Responsible AI Decision Making in Finance. The authors thank Efthymios Tsaprazlis and Sabyasachee Baruah for helpful comments.
Appendix A More Implementation Details
@@
For the baseline, we sample the random labels for the pair similarly to the random labels in LiaHR. In the demonstrations, we use equal amounts of pairs with gold labels and random labels. For Demux, we use the same training regime as Chochlakis et al. (2023), using the intra loss with a coefficient of $0.20.2$, but training only on the train set instead of integrating the dev set in training after early stopping. Standard deviation is shown for 3 model runs.
@@
Appendix B Full Human Evaluations
In human evaluations, to avoid biasing annotators towards specific answers — for example, having the dataset label always as the first option Turpin et al. (2024) —, we randomly interleave reasonable and unreasonable examples (label according to the model) for the Reasonableness check, whereas for the Preference check, we randomly change the order with which the label in the dataset and the model’s prediction are presented to the annotators. We present full results, including the number of trials and the precise numbers of them, here in Table 4 and Table 5. We note that we asked 6 colleagues (all male, ages 20-28, student researchers) to annotate to get as many perspectives as possible and avoid biasing the result. Note that the annotators were shown the Reasonableness baseline prompt from Table 3, modified appropriately.
@@ Appendix B Full Human Evaluations
Appendix C More models on SemEval properties
@@
Appendix D MFRC properties
@@
Appendix E Results on objective tasks
@@
Appendix F Degradation in copy-paste performance
@@
[…]
@@
Appendix G Extra Ecological Validity results
@@
Appendix H Filtering per Label
@@
[…]
@@
[…]
@@
[…]
Table 9: Success rates of LiaHR on GoEmotions using 25-shot GPT-4o.
Appendix I Position of Label in the Haystack
We also experiment with changing the position of the query in the prompt and evaluating how all our metrics change. We present our results in Figure 14. We see, remarkably, that no major changes are observed in the predictions of the model, irrespective of where the query appears in the demonstrations. It is very interesting to see that even when the query is the last demonstration (just before itself then), the results remain remarkably similar to when it appears first in the prompt, separated by 15 examples with itself.

### v2

@@ Acknowledgments
This project was supported in part by funds from DARPA under contract HR001121C0168, NSF CIVIC, and USC-Capital One Center for Responsible AI Decision Making in Finance. The authors thank Efthymios Tsaprazlis, Efthymios Georgiou, Kleanthis Avramidis and Sabyasachee Baruah for helpful comments.
Appendix A Methodology
In this section, we present a mathematical formulation for the baseline and LiaHR to avoid any potential ambiguities arising from the natural language description of the main text. We follow the notation of Chochlakis et al. (2025). For a set of examples $\mathcal{X}$, and a set of labels $\mathcal{Y}$, a dataset $\mathcal{D}^{a}$ defines a mapping $f^{a}:\mathcal{X}\rightarrow\mathcal{Y}$, where $a$ denotes a specific annotator or the aggregate. Similarly, $\mathcal{D}^{a}=\{(x,y)\ |\ x\in\mathcal{X},y=f^{a}(x)\}$, which is characterized by joint distribution $p^{a}(x,y)$. The gold query pair is denoted as $(x_{q},y^{a}_{q}),\ y^{a}_{q}=f^{a}(x_{q})$.
Reasonableness Baseline
We want to sample $k$ train documents from $\mathcal{X}$ to create a prompt with document-label pairs, as well as corresponding binary reasonableness labels, denoted simply as $1$ and 0666in the actual prompt, the labels are presented as “yes” or “no”, or “reasonable” and “unreasonable”, not as $1$ and 0, as shown in Table 3.. We choose half ($\frac{k}{2}$) pairs to have the “reasonable” label, and for other half the “unreasonable” label. To sample reasonable pairs for our prompt, we sample document-label pairs directly from $\mathcal{D}^{a}$ as $S^{r}=\{(x_{i},y_{i},1):\ (x_{i},y_{i})\sim p^{a},\ i\in[\frac{k}{2}]\}$. For unreasonable pairs, we sample the documents $x$ and the labels $y$ independently of each other from the dataset as $S^{u}=\{(x_{i},y_{i},0):\ (x_{i},y_{i})\sim p^{a}_{I},\ i\in[\frac{k}{2}]\}$, where $p^{a}_{I}(x,y)=p^{a}(x)p^{a}(y)$, in effect assigning random yet in-distribution labels to each document. The complete demonstrations for the model are $S=S^{r}\cup S^{u}$, and the order of the examples in the prompt is determined randomly. The query document is presented with the gold label $y_{d}^{a}=y^{a}_{q}$ (like $S^{r}$; “Gold pair” in the baseline figures like Figure 3) or a random label $y_{d}^{a}$ independently from the query $x_{q}$ using $p^{a}(y)$ (like $S^{u}$; “Rand pair” in the baseline figures) at the end, and we elicit the final prediction from the model, namely $1$ or 0: $S^{\prime}=S\cup\{(x_{q},y_{d}^{a},\_)\}$. We show the rate of $1$ predictions in our results.
LiaHR
To create the prompt, we sample $k$ demonstrations for the prompt from $\mathcal{D}^{a}$ with $p^{a}$ as $S=\{(x_{i},y_{i}):\ (x_{i},y_{i})\sim p^{a},\ i\in[k]\}$. In LiaHR, the first demonstration is the query $x_{q}$. For this demonstration, we either use its gold label $y_{d}^{a}=y_{q}^{a}$ (“Query w/ gold” in the LiaHR figures like Figure 2) or sample a random label $y_{d}^{a}$ independently from the query $x_{q}$ using $p^{a}(y)$ (“Query w/ rand” in the LiaHR figures). This query pair is included in the prompt as the first demonstration, and the query document $x_{q}$ is appended in the prompt, eliciting a prediction for it from the model, $S^{\prime}=\{(x_{q},y_{d}^{a})\}\cup S\cup\{(x_{q},\_)\}$. We measure and show the similarity between the predictions of the model with $y_{d}^{a}$, and only measure the similarity of the prediction with $y_{q}^{a}$ for “Gold perf w/ rand” in LiaHR figures.
Appendix B More Implementation Details
@@
For the baseline, we sample the random labels for the pair similarly to the random labels in LiaHR. In the demonstrations, we use equal amounts of pairs with gold labels and random labels. For Demux, we use the same training regime as Chochlakis et al. (2023), using the intra loss with a coefficient of $0.2$, but training only on the train set instead of integrating the dev set in training after early stopping. Confidence intervals are shown from 3 model runs.
@@
Appendix C Full Human Evaluations
In human evaluations, to avoid biasing annotators towards specific answers — for example, having the dataset label always as the first option Turpin et al. (2024) —, we randomly interleave reasonable and unreasonable examples (label according to the model) for the Reasonableness check, whereas for the Preference check, we randomly change the order with which the label in the dataset and the model’s prediction are presented to the annotators. We present full results, including the number of trials and the precise numbers of them, here in Table 4 and Table 5. We note that we recruited 11 colleagues (6 males and 5 females, ages 20-28, students or researchers) to annotate to get as many perspectives as possible and avoid biasing the result. Note that the annotators were shown the Reasonableness baseline prompt from Table 3, modified appropriately.
@@ Appendix C Full Human Evaluations
Appendix D More models on SemEval properties
@@
Appendix E MFRC properties
@@
Appendix F Results on objective tasks
@@
Appendix G Degradation in copy-paste performance
@@
Original
$0.623{\scriptstyle\pm 0.001}$
$\textbf{0.574}{\scriptstyle\pm 0.001}$
Replaced
$\textbf{0.624}{\scriptstyle\pm 0.002}$
$\textbf{0.574}{\scriptstyle\pm 0.003}$
Replaced (trn)
$0.615{\scriptstyle\pm 0.001}$
$0.562{\scriptstyle\pm 0.001}$
Filtered
$\textbf{0.624}{\scriptstyle\pm 0.003}$
$0.561{\scriptstyle\pm 0.002}$
Bsl Filtered
$0.615{\scriptstyle\pm 0.002}$
$0.558{\scriptstyle\pm 0.002}$
Predictions
$0.430{\scriptstyle\pm 0.004}$
$0.474{\scriptstyle\pm 0.000}$
@@
Appendix H Extra Ecological Validity results
@@
Appendix I Filtering per Label
@@
anger
$0.972{\scriptstyle\pm 0.016}$
anticipation
$0.921{\scriptstyle\pm 0.017}$
disgust
$0.939{\scriptstyle\pm 0.019}$
fear
$0.977{\scriptstyle\pm 0.016}$
joy
$0.965{\scriptstyle\pm 0.010}$
love
$0.973{\scriptstyle\pm 0.019}$
optimism
$0.995{\scriptstyle\pm 0.007}$
pessimism
$0.922{\scriptstyle\pm 0.034}$
sadness
$0.994{\scriptstyle\pm 0.008}$
surprise
$1.000{\scriptstyle\pm 0.000}$
trust
$0.867{\scriptstyle\pm 0.094}$
@@
admiration
$0.950{\scriptstyle\pm 0.021}$
anger
$0.973{\scriptstyle\pm 0.000}$
fear
$1.000{\scriptstyle\pm 0.000}$
joy
$0.871{\scriptstyle\pm 0.020}$
optimism
$0.908{\scriptstyle\pm 0.036}$
sadness
$0.930{\scriptstyle\pm 0.028}$
surprise
$0.944{\scriptstyle\pm 0.020}$
@@
authority
$0.889{\scriptstyle\pm 0.157}$
care
$0.939{\scriptstyle\pm 0.043}$
equality
$0.978{\scriptstyle\pm 0.031}$
loyalty
$0.974{\scriptstyle\pm 0.036}$
proportionality
$1.000{\scriptstyle\pm 0.000}$
Table 9: Success rates of LiaHR on GoEmotions using 25-shot GPT-4o.
Appendix J Position of Label in the Haystack
We also experiment with changing the position of the query in the prompt and evaluating how all our metrics change. We present our results in Figure 14, with standard deviations shown. We see that no major changes are observed in the predictions of the model, irrespective of where the query appears in the demonstrations. It is very interesting to see that even when the query is the last demonstration (just before itself then), the results remain remarkably similar to when it appears first in the prompt, separated by 15 examples with itself.
Appendix K Overall Reasonableness of Annotations
@@
$\begin{split}p(\text{reasonable})&=\\ p(\text{reasonable}\ &|\ \text{{\mbox{LiaHR}} reasonable})\\ &\cdot p(\text{{\mbox{LiaHR}} reasonable})\\ +p(\text{reasonable}\ &|\ \text{{\mbox{LiaHR}} unreasonable})\\ &\cdot(1-p(\text{{\mbox{LiaHR}} reasonable}))\\ &=0.682.\end{split}$ (1)
The same estimate, when checking with Llama-3 70b, comes to $0.625$, and with GPT-3.5 to $0.727$. The results are only an approximation, since LiaHR results are presented in Jaccard Score, not accuracy. The same procedure can be used with the baseline, deriving more theoretically sound estimates.
SemEval GoEmotions MFRC
FilterBERT
   Gold pairs 0.824 $\pm$ 0.017 0.751 $\pm$ 0.010 0.181 $\pm$ 0.006
   Rand pairs 0.244 $\pm$ 0.010 0.215 $\pm$ 0.011 0.039 $\pm$ 0.008
GPT-4o (5-shot)
   Gold pairs 0.887 $\pm$ 0.078 0.693 $\pm$ 0.063 0.430 $\pm$ 0.254
   Rand pairs 0.277 $\pm$ 0.235 0.243 $\pm$ 0.029 0.043 $\pm$ 0.038
Table 10: Filtering accuracy for “reasonable” vs. “unreasonable” label-document pairs using a proxy supervised BERT-based classifier (trained on all data) and GPT-4o (5-shot). Gold pairs match the true label; Rand pairs use randomly sampled labels.
Appendix L BERT-based Filtering Baseline
In this section, we present results for a BERT-based filtering baseline, FilterBERT. FilterBERT’s output is the binary decision of whether to filter a document-label pair out of the data pool. It is a proxy supervised baseline, as we do not use actual annotated data for reasonableness, but instead use the same strategy as for the LLM baseline to construct data. Namely, documents that are paired with their gold label from the dataset are considered “reasonable” pairs. To create “unreasonable” pairs for a document, we sample the labels from a random document in the dataset. Practically, we use all pairs from the original dataset, and for each document we also create an “unreasonable” pair, doubling the size of the dataset. The input is formatted similarly to Demux Chochlakis et al. (2023), where the input consists of the CLS token, followed by the candidate labels, in turn followed by a SEP token, and finally the input document. An example input, therefore, is “[CLS] anger, anticipation, optimism [SEP] I DIDN’T ASK FOR THIS EITHER IT JUST HAPPENED”. We use the contextual embedding of CLS with a two-layer neural network (again, similar to Demux) to make the final binary prediction with a threshold of $0.5$ on the output sigmoid function. Training details are otherwise identical to Demux (note that we have removed the intro loss coefficient because we do not apply the classifier on each emotion of the prompt). Our results are presented in Table 10, in comparison to the 5-shot GPT-4o baseline results we have already presented in Figures 3, 5 and 9.
The BERT-based model cannot be used to conduct the ecological validity tests due to the fact that it itself needs to use the train and dev sets to be trained, so a more direct comparison is not possible with our current setting, a shortcoming of using a BERT-based model for filtering. However, from these existing results, GPT-4o seems to align more with our intuitions of how a model should perform. For SemEval, the performance of GPT-4o is closer to random baseline performance compared to BERT in rejecting emotions, and accepts more labels. For GoEmotions, the BERT-based model seems to learn the noise in GoEmotion arising from the hierarchical clustering that we apply, achieving higher acceptance rates (as we have mentioned before). The superiority of GPT-4o on MFRC is evident.